how many 2/3 acres lots can fill a 9 3/4 acres of land?
Answer by jim_thompson5910(35256)

Let x = # of lots

If each lot is 2/3 of an acre, then x lots will be total of (2/3)x acres

We want this to be 9 3/4, so...

(2/3)x = 9 3/4

(2/3)x = 9 + 3/4

(2/3)x = 36/4 + 3/4

(2/3)x = 39/4

2x = 3*(39/4)

2x = 117/4

x = (117/4)/2

x = (117/4)*(1/2)

x = 117/8

x = (112 + 5)/8

x = 112/8 + 5/8

x = 14 + 5/8

x = 14 5/8


So you need 14 5/8 lots, but you can only have a whole number of lots. So round down to get the final answer of 14 lots.